OpenStudy (anonymous):

how do you write this symbol expression four squared divided by two to the fourth power

OpenStudy (unklerhaukus):

\[4^2/2^4\]

OpenStudy (anonymous):

then how would you evaluate it

OpenStudy (anonymous):

16/16=0

OpenStudy (unklerhaukus):

\[4^2/2^4=(4\times4) / (2\times2\times2\times2)\]\[\quad\qquad=16/16\]\[\quad\qquad=\]

OpenStudy (unklerhaukus):

not zero

OpenStudy (anonymous):

1

OpenStudy (unklerhaukus):

thats it

OpenStudy (anonymous):

thanks